OpenStudy (anonymous):

Simplify: (sin Θ − cos Θ)2 + (sin Θ + cos Θ)2

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@suyogsomavanshi

OpenStudy (nurali):

(sinθ+cosθ)²+(sinθ-cosθ)²=0 (sin²θ + 2sinθcosθ + cos²θ)+(sin²θ - 2sinθcosθ + cos²θ)=0 (sin²θ+ cos²θ) + 2sinθcosθ - 2sinθcosθ +(sin²θ+ cos²θ)=0 1 + 0 +1=0 2≠0 no solutions